What is the pH of a 2.9 x 10-5 M solution of HBr?
Answer #1

Given:

[HBr] = 2.9*10^-5 M

So,

[H+] = 2.9*10^-5 M

use:

pH = -log [H+]

= -log (2.9*10^-5)

= 4.54

Answer: 4.54